at what Kekvin temperature will 37.5 mol of Ar occupy a volume of 725 L at a pressure of 675 torr ?


1
Expert's answer
2022-03-31T04:08:02-0400

At STP ie 760 torr and 273K, 1 mol occupy 22.4L.

Therefore, 37.5 mol occupy 22.4×37.5= 840L

P1= 760 torr, V1= 840L and T1= 273K

P2= 675 torr, V2= 725L and T2= ?

P1V1÷T1 = P2V2÷T2

T2= P2V2T1 ÷(P1V1)

= (675×725×273)÷(760×840)

= 209.3K